# Enhancement of Electric Drive in Silicon Quantum Dots with Electric Quadrupole Spin Resonance

### v2

@@ I Introduction
Qubits, the fundamental building block of quantum computers, work by storing quantum information in two quantum states. Therefore, implementations of a quantum computer are possible as long as there are two quantum states (for each qubit) that can be coherently controlled and manipulated to interact with other qubits. Such implementations include electron spin qubits implemented using silicon Metal-Oxide Semiconductor (MOS) technology, which offer numerous advantages. High-fidelity quantum operations have been demonstrated in silicon [1, 2, 3, 4, 5, 6] but the main advantage over other qubit technologies is the ability to leverage existing MOS technology. That allows us to build a quantum processor with up to billions of qubits that will be needed to create a universal quantum computer, thus solving problems that would otherwise be too costly for classical computers [7, 8]. The large number of qubits allows quantum error correction algorithms to be implemented, which protect fragile qubits from external sources of noise that can destroy quantum information [9].
@@
Electron spins can be coherently controlled using a microwave source to generate an AC magnetic field resonant with the electronic energy level splitting due to the Zeeman effect [7]. However, using the magnetic field for individual qubit addressability can be problematic because the field is not localized in space. There are current research efforts to address this problem by using an “always on” global magnetic field [10, 11, 12] such as the recently proposed SMART protocol [13, 14, 15, 16].
@@
Alternatively, electric driving offers another solution to this problem because electric fields are much more localized. The electron, however, requires a position and/or momentum-dependent magnetic field to couple the spin degree of freedom to the electric field. This coupling is commonly called spin orbit coupling (SOC) and can be realized with an external micro-magnet [17, 18, 19, 20], or with intrinsic SOC in silicon MOS heterostructures [21, 22], enabling Rabi frequencies up to tens of megahertz [1], which can be an order of magnitude faster compared to the typical magnetic electron spin resonance (ESR) drive [23]. There have also been novel schemes of electric driving to enhance its effect [24], as well as in-depth studies of achieving high-fidelity control [25]. However, one key advantage of using intrinsic SOC over the micromagnet is the reduction of engineering challenges in scale up [7]. The most common way to electrically drive a qubit is called Electric Dipole Spin Resonance (EDSR), where Rabi frequencies are promoted by an alternating electric field coupled with the electron orbital degree of freedom in the dipolar approximation.
@@
In a demonstration of electrically driven multielectron spin qubits using MOS quantum dots [26], it was shown that a multielectron dot not only mitigates charge noise and disorder [27], but also allows for an order of magnitude enhancement in Rabi frequency compared to a single electron [28]. In particular, with 5 and 13 electrons, a higher quality factor was observed. This is a significant step towards having fault-tolerant qubits, and therefore, we seek to understand and reliably reproduce this behavior in a scale-up architecture.
@@
In this paper, we develop a theoretical model based on the experimental results obtained in Ref. [26]. We begin with a model describing the electric driving of orbital states which controls the spin-orbital character of driven states via the electrostatic gates [22]. Here, we add to the model by arguing that the EDSR description is not sufficient to describe the observed enhancements in Rabi frequency of multielectron dots with > 3 electrons, and we propose a new method named “Electric Quadrupole Spin Resonance” (EQSR).
In the following sections, we first provide a model of the device reported in Ref. [26], making explicit the states that form the spin-orbit qubit for the 5 electrons case and how that leads to EQSR coupling. Then, we present theoretical evidence for EQSR by comparing the theoretical results of this modeling with the experimental results. Finally, we conclude by discussing how we can leverage EQSR in the future.
@@ II.1 Microscopic Model
In this section, we begin by building an effective Hamiltonian that describes the system reported in Ref. [26]. We will model the dot with an effective single-particle 2D Hamiltonian that describes the low energy physics as follows:
$H_{\mathrm{DC}}=H_{\mathrm{xy}}+H_{\mathrm{Zee}}+H_{\mathrm{SO}}\;.$ (1)
The name DC refers to the static parts of the physics - later we will look at the AC driving terms. This is a simplified description of the system, that illustrates the core physics presented here. The complete Hamiltonian is shown in the Supplementary Material [29] with additional terms relevant to the system that are necessary for fitting but not for the EQSR description.
@@
where $m_{\mathrm{t}}$ is the transverse effective mass in the $\Delta$ valley of silicon, $\delta\equiv\omega_{x}/\omega_{y}$ describes the ellipticity of the parabolic confinement ($\delta=1$ corresponds to a perfectly circular dot) and $p_{x(y)}$ is the kinetic momentum in the x(y) direction. The spin degree of freedom enters via the Zeeman Hamiltonian $H_{\mathrm{Zee}}=\frac{g\mu_{B}}{2}\mathbf{B}_{0}\cdot\bm{\sigma}$, with $\mathbf{B}_{0}=-B_{0}\hat{y}$ being a static uniform magnetic field. These two components can be treated as the unperturbed part of the Hamiltonian, and their eigenstates written as $|n,s\rangle$, where $n\in\{S,P_{x},P_{y},\dots\}$ denotes the 2D quantum harmonic oscillator orbital and $s\in\{\uparrow,\downarrow\}$ denotes the spin state.
The spin-orbit component $H_{\mathrm{SO}}$ is considered as a perturbation to $H_{\mathrm{xy}}+H_{\mathrm{Zee}}$. It promotes a hybridization of orbital and spin degrees of freedom between the unperturbed eigenstates $|n,s\rangle$. In this analysis, it can be written as
@@
where $h_{\mathrm{R}},h_{\mathrm{D}}$ are the Rashba and Dresselhaus terms with respective coefficients $\alpha,\beta$ which arise due to the lowered symmetry of the heterostructure and Si/SiO2 interface respectively [30]. $\sigma^{\prime}$ indicates the Pauli matrices in the spin coordinates as opposed to the real space coordinates and crystallographic axes. The relations between the different coordinates in the system are described in the Supplementary Material [29] and are consistent with the definition of $H_{\mathrm{Zee}}$.
Figure 1: Schematic of p-orbital-like states. As the ellipticity of the electrostatic potential decreases, SOC hybridization of the orbital and spin characters of the states increases, forming a spin-orbit qubit.
In the 5-electron experiment, the qubit is defined by the unperturbed states $|0\rangle_{\mathrm{unp}}=|P_{y},\uparrow\rangle$ and $|1\rangle_{\mathrm{unp}}=|P_{y},\downarrow\rangle$. However, $H_{\mathrm{SO}}$ couples the different orbital states, resulting in these states becoming,
@@
where $a_{0}$ comes from first order effect of $H_{\mathrm{SO}}$ (from the transition $|{P_{y},\uparrow}\rangle\xrightarrow{k_{y}\sigma_{x}^{\prime}}|{S,\downarrow}\rangle$), $b_{0}$ comes from second order effect of $H_{\mathrm{SO}}$ (from the transition $|{P_{y},\uparrow}\rangle\xrightarrow{k_{y}\sigma_{x}^{\prime}}|{S,\downarrow}\rangle\xrightarrow{k_{x}\sigma_{z}^{\prime}}|{P_{x},\downarrow}\rangle$), and $\mathcal{N}_{0}$ is a normalization constant. In a similar fashion the coefficients $a_{1}$, $b_{1}$ and $\mathcal{N}_{1}$ are defined. The approximation on the right hand side of the equation indicates that the original, unperturbed states are still the most dominant component. The states in Eq. (4) differ both by spin and orbit degrees of freedom and the hybridization promoted by the spin-orbit Hamiltonian is increased when the dot ellipticity decreases (that is, when $\delta$ approaches 1), as we will analyze in the next section. Fig. 1 shows a schematic of this hybridization for p-orbital states.
Based on Eq. (3) and the coupling pathways of the states via $H_{\mathrm{SO}}$, we can determine that the $a_{0(1)}$ coefficients are proportional to $\alpha+\beta$ and the $b_{0(1)}$ coefficients are proportional to $\alpha^{2}-\beta^{2}$. These proportionalities lead to different dependencies of the Rabi speed-up on spin-orbit coupling.
@@ II.2 Electric Quadrupole Spin Resonance (EQSR)
Now that we have described the DC component of the Hamiltonian, we will describe the AC driving component. An alternating electric field $\mathbf{E}=E_{y}\cos(\omega t)\hat{y}$ interacts with an electron according to the following Hamiltonian:
@@
This method of driving is called Electric Dipole Spin Resonance (EDSR). With 5 electrons present in the dot, the states $|0\rangle$ and $|1\rangle$ are those described in Eq. (4), respectively. Therefore, the Rabi frequency would depend on the $a_{0(1)}$ coefficients, since the $\hat{y}$ operator connects states differing in one orbital excitation in the $y$ direction, in this case, the states $|P_{y}\rangle$ and $|S\rangle$. This means that a Rabi frequency enhancement would only occur when $a_{0(1)}$ are large, i.e. when the $|P_{y}\rangle$ and $|S\rangle$ orbitals come close in energy, enabling spin-orbit coupling to hybridize the two in a significant fraction. However, the large measured orbital splitting of $\sim 4~{}\mathrm{meV}$ shows that this cannot be the case, indicating that EDSR cannot be responsible for the observed Rabi enhancement.
@@
Then, what can cause the enhancement? The experiments give us two clues. First, no enhancement is observed for a single electron dot [26], suggesting that the enhancement might depend on a property of the $P$ or $D$ orbitals, not present in the $S$ orbitals. The main difference here is the 2- or 3-fold orbital degeneracy.
Second, as the multi-electron dot is tuned towards the degeneracy point (by increasing the electrostatic gate voltage $V_{\mathrm{G2}}$), the lateral electrostatic confinement is altered. We confirm this via real-space simulations of the COMSOL electrostatic simulations of the device at the operating point, extracting only the single-particle states from electronic structure simulation tools [31]. From the simulations, we obtained the excitation energies $\mathcal{E}_{x}=\hbar\omega_{x}$ (circles) and $\mathcal{E}_{y}=\hbar\omega_{y}$ (diamonds) in the 5 (blue) and 13 (red) electron regimes, plotted in Fig. 2(a). All the energy traces appear to be linear with the gate voltage. We also plot in Fig. 2(b) the extracted ellipticity, $\delta$, by taking the ratio between $\mathcal{E}_{x}$ and $\mathcal{E}_{y}$ in both the 5 and 13 electron configurations (blue and red respectively). We observe that $\delta$ varies linearly as a function of gate voltage $V_{\mathrm{G2}}$. This is also further confirmation that no significant Rabi enhancement can occur in the electric dipole spin resonance regime, since $\omega_{y}$ does not vary significantly with the shape of the dot (and neither does $\omega_{x}$ vary enough to cause this enhancement), given that $\Omega\propto y$ and $y\propto\frac{1}{\sqrt{\omega_{y}}}$. Furthermore, from Fig. 2(b), it is found that the degeneracy point is only achieved when the dot becomes more circular ($\delta$ approaching unity as excitation energies decrease), which is opposite to the expected electric dipole effect.
In order to account for these observations in Fig. 2, our model should include the second order Taylor expansion of the driving electric field, or in other words, the quadrupole field
@@
Here, we see that the Rabi frequency enhancement can also be promoted by the second-order perturbation terms of the states $|0\rangle$ and $|1\rangle$. In fact, as the energy difference between the $P_{x}$ and $P_{y}$ orbitals approaches that of the Zeeman splitting, which can be achieved by tuning the gate voltages, the coefficient $|b_{i}|$ grows larger. Therefore, this treatment may predict the enhancement observed within this regime. For the 13-electron experiment, this quadrupole driving will also couple the $D_{xy}$ with $D_{yy}$ or $D_{xx}$ orbitals.
Although quadrupole coupling is expected due to the non-uniformity of the electric field produced by the gates, it is often been neglected in calculations of the Rabi frequency. The $Q_{xy}$ term in the Hamiltonian is the only term to directly couple the $P_{x},P_{y}$ orbitals, as well as the $D_{xy},D_{yy}$ ones, which can account for the observed Rabi enhancement. There are other components of the quadrupole field in the driving term $H_{\mathrm{AC}}$, but they do not offer direct coupling between the orbitals and therefore become negligible in our description.
We will conclude this section with some intuition on the influence of the quadrupole $Q_{xy}$ term on the dot. Whereas electric dipole spin resonance causes a lateral oscillation of the dot, EQSR causes a dilation/contraction of potential along the $x=\pm y$ axes (see Fig. 3). Another way of thinking about this driving term is that it has an $x=y$ symmetry axis, which allows it to couple orbitals of opposite $x$ and $y$ symmetry. COMSOL simulations yield $Q_{xy}/E_{y}=1.2\times 10^{-3}$ nm-1, a ratio that we kept fixed when fitting our model to the experiment. Our dot occupies a region of linear size $L\sim 20$ nm, and therefore we see that $Q_{xy}L/E_{y}\ll 1$, as required for the validity of the multipolar expansion.

### v3

@@ I Introduction
Qubits are the fundamental building blocks of quantum computers and must be coherently controlled and manipulated to interact with other qubits. Such implementations include electron spin qubits implemented using silicon Metal-Oxide Semiconductor (MOS) technology, which offer numerous advantages. High-fidelity quantum operations have been demonstrated in silicon [1, 2, 3, 4, 5, 6] but the main advantage over other qubit technologies is the ability to leverage existing MOS technology. That provides the possibility of creating a universal quantum computer that could solve problems otherwise too costly for classical computers [7, 8]. The large number of qubits allows quantum error correction algorithms to be implemented, protecting fragile qubits from external sources of noise that can destroy quantum information [9].
@@
Electron spins can be coherently controlled using a microwave source to generate an AC magnetic field resonant with the electronic energy level splitting due to the Zeeman effect [7]. However, using the magnetic field for individual qubit addressability can be challenging because the field is not localized in space.
@@
Alternatively, electric driving is localized but the electron requires a position and/or momentum-dependent magnetic field to couple the spin degree of freedom to the electric field. This spin-orbit coupling (SOC) can be realized with an external micro-magnet [10, 11, 12, 13], or with intrinsic SOC in silicon MOS heterostructures [14, 15], enabling Rabi frequencies up to tens of megahertz [1], which can be an order of magnitude faster compared to the typical magnetic electron spin resonance (ESR) drive [16]. The most common way to electrically drive a qubit is called Electric Dipole Spin Resonance (EDSR), where Rabi frequencies are promoted by an alternating electric field coupled with the electron orbital degree of freedom in the dipolar approximation.
@@
In a demonstration of electrically driven multielectron spin qubits using MOS quantum dots [17], it was shown that a multielectron dot not only mitigates charge noise and disorder [18], but also allows for an order of magnitude enhancement that does not occur for a single electron [19]. In particular, with 5 and 13 electrons, a higher quality factor was observed. This is significant and we seek to understand and reliably reproduce this behavior in a scale-up architecture.
@@
In this paper, we develop a theoretical model based on the experimental results obtained in Ref. [17]. We begin with a model describing the electric driving of orbital states which controls the spin-orbital character of driven states via the electrostatic gates [15]. Then, we add to the model by arguing that the EDSR description is not sufficient to describe the observed enhancements in Rabi frequency of multielectron dots with > 3 electrons, and we propose a new method named “Electric Quadrupole Spin Resonance” (EQSR).
@@ II.1 Microscopic Model
In this section, we build the effective Hamiltonian that describes the system reported in Ref. [17]. We will model the dot with an effective single-particle 2D Hamiltonian
$H_{\mathrm{DC}}=H_{\mathrm{xy}}+H_{\mathrm{Zee}}+H_{\mathrm{V}}+H_{\mathrm{SO}}\;.$ (1)
The name DC refers to the static parts of the physics - later we will look at the AC driving terms. This simplified Hamiltonian contains the essential physics, for the complete one used in the fitting see[20].
@@
where $m_{\mathrm{t}}$ is the transverse effective mass in the $\Delta$ valley of silicon, $\delta\equiv\omega_{x}/\omega_{y}$ describes the ellipticity of the parabolic confinement ($\delta=1$ corresponds to a perfectly circular dot) and $p_{x(y)}$ is the kinetic momentum in the x(y) direction. The spin degree of freedom enters via the Zeeman Hamiltonian $H_{\mathrm{Zee}}=\frac{g\mu_{B}}{2}\mathbf{B}_{0}\cdot\bm{\sigma}$, with $\mathbf{B}_{0}$ being a static uniform magnetic field. $H_{\mathrm{V}}$ describes the degeneracy split of two z-valley states, due to the Si/SiO2 interface, with an energy splitting likely between 500500 $\mathrm{\mu eV}$ to $1$ meV [17, 21]. These three components can be treated as the unperturbed part of the Hamiltonian, and their eigenstates written as $|n,s,v_{\pm}\rangle$, where $n\in\{S,P_{x},P_{y},\dots\}$ denotes the 2D quantum harmonic oscillator orbital, $s\in\{\uparrow,\downarrow\}$ denotes the spin state and $v_{\pm}$ denotes the two valley states. The valley degree of freedom will not be relevant in the qubit dynamics (see below), so from now on we omit it.
The spin-orbit component $H_{\mathrm{SO}}$ is considered as a perturbation to $H_{\mathrm{xy}}+H_{\mathrm{Zee}}$. It promotes a hybridization of orbital and spin degrees of freedom between the unperturbed eigenstates $|n,s\rangle$. Here we write
@@
where $h_{\mathrm{R}},h_{\mathrm{D}}$ are the Rashba and Dresselhaus terms with respective coefficients $\alpha,\beta$ which arise due to the lowered symmetry of the heterostructure and Si/SiO2 interface respectively [22]. $\sigma^{\prime}$ indicates the Pauli matrices in the spin coordinates [20].
Figure 1: Schematic of p-orbital-like states. As the ellipticity of the electrostatic potential decreases, SOC hybridization of the orbital and spin characters of the states increases, forming a spin-orbit qubit. $\mathcal{E}_{\text{Zee}}$ is the Zeeman splitting.
The experiments of Ref. [17] were done with 5 and 13 electrons, which can be approximated within single-particle theory for the valence electron because 4 and 12 electrons are closed shell configurations.
In the 5-electron experiment, the qubit is defined by the unperturbed states $|0\rangle_{\mathrm{unp}}=|P_{y},\uparrow\rangle$ and $|1\rangle_{\mathrm{unp}}=|P_{y},\downarrow\rangle$. However, $H_{\mathrm{SO}}$ couples the different orbital states, resulting in these states becoming
@@
where $a_{0}$ comes from first order effect of $H_{\mathrm{SO}}$ (from the transition $|{P_{y},\uparrow}\rangle\xrightarrow{k_{y}\sigma_{x}^{\prime}}|{S,\downarrow}\rangle$), $b_{0}$ comes from second order effect of $H_{\mathrm{SO}}$ (from the transition $|{P_{y},\uparrow}\rangle\xrightarrow{k_{y}\sigma_{x}^{\prime}}|{S,\downarrow}\rangle\xrightarrow{k_{x}\sigma_{z}^{\prime}}|{P_{x},\downarrow}\rangle$), and $\mathcal{N}_{0}$ is a normalization constant. In a similar fashion the coefficients $a_{1}$, $b_{1}$ and $\mathcal{N}_{1}$ are defined. The approximation on the right hand side of the equation indicates that the original, unperturbed states are still the most dominant component. The states in Eq. (4) differ both by spin and orbit degrees of freedom and the hybridization promoted by the spin-orbit Hamiltonian is increased when the dot ellipticity decreases (that is, when $\delta$ approaches 1), as analyzed in the next section. Fig. 1 shows a schematic of this hybridization for p-orbital states. The hybridization of spin and orbit degrees of freedom allows to drive the qubit electrically, while the valley,for our purposes, means only that we can place two electrons per $|n,s\rangle$ state (justifying its omission from now on).
Based on Eq. (3) and the coupling pathways of the states via $H_{\mathrm{SO}}$, we can determine that $a_{0(1)}\propto\alpha+\beta$ and $b_{0(1)}\propto\alpha^{2}-\beta^{2}$ [20]. These proportionalities lead to different dependencies of the Rabi speed-up on spin-orbit coupling.
@@ II.2 Electric Quadrupole Spin Resonance (EQSR)
Let us now describe the AC driving component. An alternating electric field $\mathbf{E}=E_{y}\cos(\omega t)\hat{y}$ interacts with an electron according to the following Hamiltonian:
@@
This method of driving is called Electric Dipole Spin Resonance (EDSR). With 5 electrons present in the dot, the states $|0\rangle$ and $|1\rangle$ are those described in Eq. (4), respectively. Therefore, the Rabi frequency would depend on the $a_{0(1)}$ coefficients, since the $\hat{y}$ operator connects states differing in one orbital excitation in the $y$ direction, in this case, the states $|P_{y}\rangle$ and $|S\rangle$. This means that a Rabi frequency enhancement would only occur when $a_{0}$ is large, i.e. when the $|P_{y},\uparrow\rangle$ and $|S,\downarrow\rangle$ orbitals come close in energy: $\hbar\omega_{y}\sim\mathcal{E}_{\text{Zee}}$, enabling spin-orbit coupling to hybridize the two in a significant fraction ($\mathcal{E}_{\text{Zee}}$ is the Zeeman energy splitting). However, for typical values used in experiments, $\hbar\omega_{y}$ ($\sim 10\mathrm{meV}$ [23]) is many orders of magnitude larger than $\mathcal{E}_{\text{Zee}}$ ($\sim 100\mathrm{\mu eV}$ [17]). This indicates that EDSR cannot be responsible for the observed Rabi enhancement.
@@
Then, what can cause the enhancement? The experiments give us two clues. First, no enhancement is observed for a single electron dot [17], suggesting a possible dependence on a property of the $P$ or $D$ orbitals, not present in the $S$ orbitals. The main difference here is the 2- or 3-fold orbital degeneracy.
Second, as the multi-electron dot is tuned towards the degeneracy point (by increasing the electrostatic gate voltage $V_{\mathrm{G2}}$), the lateral electrostatic confinement is altered. We confirm this via real-space simulations of the COMSOL electrostatic simulations of the device at the operating point, extracting only the single-particle states from electronic structure simulation tools [24]. From the simulations, we obtained the excitation energies $\mathcal{E}_{x}=\hbar\omega_{x}$ (circles) and $\mathcal{E}_{y}=\hbar\omega_{y}$ (diamonds) in the 5 (blue) and 13 (red) electron regimes, plotted in Fig. 2(a). All the energy traces appear to be linear with the gate voltage. We also plot in Fig. 2(b) the extracted ellipticity, $\delta$, by taking the ratio between $\mathcal{E}_{x}$ and $\mathcal{E}_{y}$ in both the 5 and 13 electron configurations (blue and red respectively). We observe that $\delta$ varies linearly as a function of gate voltage $V_{\mathrm{G2}}$. In the dipolar coupling, the Rabi frequency given in Eq.(6) is proportional to $1/\sqrt{\omega_{y}}$ (present in the $y$ component of the dipole operator transition element). From Fig. 2(a), $\omega_{y}$ variation is very small, ruling out the possibility that the observed Rabi enhancement occurred because of non expected changes in $\omega_{y}$ when varying $V_{\mathrm{G2}}$.
To account for these observations, our model should include the second order Taylor expansion of the driving electric field, or in other words, the quadrupole field
@@
Here, we see that the Rabi frequency enhancement can also be promoted by the second-order perturbation terms of the states $|0\rangle$ and $|1\rangle$. In fact, as the energy difference between the $P_{x}$ and $P_{y}$ orbitals approaches that of the Zeeman splitting ($\hbar|\omega_{x}-\omega_{y}|\sim\mathcal{E}_{\text{Zee}}$) which can be achieved by tuning the gate voltages, the coefficient $|b_{i}|$ grows larger. Therefore, this treatment may predict the enhancement observed within this regime. For the 13-electron experiment, this quadrupole driving will also couple the $D_{xy}$ with $D_{yy}$ or $D_{xx}$ orbitals.
Although quadrupole coupling is expected due to the non-uniformity of the electric field produced by the gates, it is often been neglected in calculations of the Rabi frequency. The $Q_{xy}$ term in the Hamiltonian is the only quadrupole component to directly couple the $P_{x},P_{y}$ orbitals, as well as the $D_{xy},D_{yy}$ ones, which can account for the observed Rabi enhancement. For this reason we neglect the other components.
To conclude this section, whereas EDSR arises from a lateral oscillation of the dot, EQSR is caused by a dilation/contraction of potential along the $x=\pm y$ axes (see Fig. 3). Another way of thinking about this driving term is that it has an $x=y$ symmetry axis, which allows it to couple orbitals of opposite $x$ and $y$ symmetry. COMSOL simulations yield $Q_{xy}/E_{y}=1.2\times 10^{-3}$ nm-1, a ratio that we kept fixed when fitting our model to the experiment. Our dot occupies a region of linear size $L\sim 20$ nm, and therefore we see that $Q_{xy}L/E_{y}\ll 1$, as required for the validity of the multipolar expansion.